Solve the system of equations and enter the solution as an ordered pair.
4x - 7y = -23
8x + 5y = 49

Answer by richwmiller(17219)

4x - 7y = -23
8x + 5y = 49
(3,5)
multiply first by 2
8x-14y=-46
8x+ 5y =49
subtract
-19y=-46-49
-19y=-95
y=5
8x+25=49
8x=24
x=3